# A Comprehensive Survey of Advanced Persistent Threat Attribution: Taxonomy, Methods, Challenges and Open Research Problems

### v1

@@ V-A Automated Attribution Framework
V-B APT Attribution Dataset Classification
We curate the publicly available datasets, which can be used to develop APT attribution methodology. In our exploration, we classify public data sources based on the type of data they consist. Therefore, we classify the data repositories into four types: Malware dataset, Threat report dataset, Attack pattern dataset, and Heterogeneous dataset. The distribution is shown in Fig. 6. We find a total of twelve public sources that facilitate data collection specific to APT attacks, which we discuss in detail below:
Figure 6: Dataset Classification of APT Attribution
@@ V-B2 Threat Report Dataset
V-B3 Attack Pattern Dataset
MITRE ATT&CK framework [118] also lists a set of TTPs used by threat groups in past attributed attack campaigns. We can map the TTPs associated with a specific threat group through ATT&CK framework [80].
V-B4 Heterogeneous Dataset
An Excel sheet focused on APT groups and their operations [103] likely is a critical asset for cybersecurity analysts and researchers. It encompasses a wealth of data, including identifying APT groups, their geographical or organizational origins, details on TTPs employed, associated malware tools, descriptions of known campaigns, and IoCs. This organized compilation facilitates a comprehensive understanding of the threat landscape posed by these sophisticated actors, detailing their methodologies, and the specific vulnerabilities they exploit. By aggregating and categorizing this information, the document aids in identifying, analyzing, and mitigating threats from various APT groups, enhancing the overall cybersecurity posture against such entities.
MITRE curates the extensive details related to activities performed by threat groups [106]. It endeavors to map APT groups to their TTPs [118], as well as the softwares [111] attackers utilize and the campaigns [115] they conduct, and various naming aliases which acknowledge that different organizations may track the same entities under different names. This approach aids in understanding the overlaps and distinctions among groups, enhancing analyst awareness and encouraging further investigation. Software in MITRE framework encompasses a variety of applications, ranging from custom malware to commonly used operating system utilities, categorized into malware and tools based on their intended use by adversaries. This categorization helps understand the specific techniques that different software instances can employ, mapped to the groups known to use them. Campaigns are defined as structured series of intrusion activities with specific objectives, targets, and duration. MITRE records these activities, providing unique identifiers for unnamed operations or using publicly reported names, which aims to attribute them to known groups and their used software where possible. MITRE facilitates a comprehensive view of threat actor’s operations over time and offers insights into the evolution of their tactics and methodologies.
Laurenza and Lazzeretti introduce “dAPTaset”  [26], a structured database for APT information. The database structure includes APTs, keywords, threat reports, samples, network information, CVEs (Common Vulnerabilities and Exposures), techniques, and more. The data collection process involves multiple sources such as MITRE ATT&CK [118], APT groups and operations spreadsheets [103], APTNotes [19], and MISP galaxy cluster [13], from which APT names, aliases, victims, operations, and external references are extracted. They parsed the public threat reports to gather IoCs, file hashes, IP addresses, network URLs, emails, CVEs, and keywords related to APTs. The dataset is further enriched by cross-referencing with cybersecurity websites like VirusTotal [16], VirusShare [27], Malshare [34], and Malpedia [47] to validate and supplement information. This dataset uses $821$ reports of $88$ different threat groups. In addition, it also includes a total of $21,841$ binary hashes, $2,620$ network-related IoCs, $8,927$ unique files, $169$ CVEs, and $175$ different techniques related to APT activities. The dataset is publicly available [91] for the benefit of the community.
V-C Classification of Attribution methods
@@
V-C1 Malware-based methods
Kida et al. [29] present a novel way to leverage malware’s fuzzy hashes to attribute APT groups. They calculated five types of fuzzy hashes: TLSH, SSDEEP, SDHASH, IMPFUZZY, and LZJD for all malware samples and used computed hash as natural language input to the classifier. After hash calculation, the hashes are encoded using the NATO encoding method, and they use the padding method to keep the encoded length the same. Further, they use character level and n-gram vector embeddings to convert encoded hashes into feature vectors. The embedding vectors are forwarded to four different ML classifiers: RF, SVM, KNN, and NB. Their result demonstrates that RF with IMPFUZZY hash performs best of all implemented methods.
Rosenberg et al. [79] propose a method named DeepAPT to employ raw dynamic malware features and attribute APT groups linked with the given malware sample. They execute samples in a sandbox environment and collect behavior reports for each malware sample. DeepAPT considers words present in behavior reports as features and evaluates the frequency of each unique word in all behavior reports. The authors selected the top $50,000$ frequent words as features and evaluated one-hot encoding to prepare the dataset. Further, they employ a 10-layer DNN to learn patterns from raw dynamic features for APT group classification.
@@ V-C1 Malware-based methods
Haddadpajouh et al. [116] introduce a multi-view fuzzy consensus clustering model to attribute cyber threat payloads, specifically malware, to their respective actors. Their study follows a multi-view approach (multiple pieces of evidence) rather than single-view (single piece of evidence) analyses. The multi-view is generated using four categories of features: Header, System call, Byte code, and Opcode. The authors calculate multi-view for each feature category using binary vector (1 if present 0 if absent), count vector (number of it appears), frequency of occurrence, TF-IDF vector, and eigenvector of CFGs. Further, they utilize fuzzy pattern trees and multi-modal fuzzy classifiers for inference. Their research demonstrates that a multi-view approach is significantly better than single-view analyses for cyber threat attribution.
@@ V-C4 Heterogeneous data-based methods
For malware-based methods, authors in literature extract the following types of features from malware samples: words from cuckoo sandbox reports, OpCode, ByteCode, API call, Header data, IR (Intermediate Representation) codes, strings, Fuzzy hashes (TLSH, SSDEEP, SDHASH, IMPFUZZY, and LZJD), CFG, and runtime modules. Threat report-based attribution methods extract words from threat intelligence reports as features. Attack pattern-based methods leverage TTPs as a feature for attribution. Meanwhile, heterogeneous data-based attribution methods leverage diverse sets of features like TTPs, software used, and heterogeneous attributed graph schemas. The distribution of feature types used in the literature is shown in Fig. 7(a).
Out of all features mentioned in the literature TTPs and threat report texts has been used widely to present attribution method. For feature transformation, the literature leverages the following different methods: One-Hot encoding, SMOBI word embedding, TF-IDF, Count, Statistical Characteristics, SIMVER word embedding, n-gram method, NATO encoding, Basic block characteristics, frequency analysis, Node embedding, BERT embedding, and Node2vec embedding. The distribution of type of transformation methods has been used in the literature is shown in Fig. 7(b). Out of all used transformation methods One-Hot encoding and TF-IDF has been used widely in the literature.
@@
(a) Feature types
(b) Feature transformation methods
(c) Algorithm types
Figure 7: Distribution of feature types (a), transformation methods (b), and algorithm types (c)
@@ V-D Discussion
(P1) As malware can be tweaked, distributed, or sold, i.e., given the rise of MaaS (Malware-as-a-Service) [33], it may lead to notable challenges in uncovering the actual origin of a cyber attack. Malwares’ behaviour-based threat attribution may provide an inconsistent view of threat actors when the same threat actor employs numerous malware in various campaigns. (P2) Malware-based threat attribution can be misleading in the case of collaborative and multi-layered attacks, as these attacks do not rely on a single instance of malware. (P3) For the threat-report-based attribution methods, authors employ threat report words as features for classification in the literature. Comparing the words of two threat reports shows the similarity between threat reports rather than the similarity between the threat actor’s behavior or modus operandi. Comparing words for document similarity has been studied widely in the natural language processing domain [28, 38]. (P4) Given P3, such methods need to reverse engineer and identify words contributing to the attribution or profiling of a specific threat actor. (P5) Attack pattern-based methods leverage TTPs in the literature. TTP of a threat group can evolve over time [52], and TTP-based threat attribution can miss-attribute in such cases. Therefore, TTP-based attribution methods demand continuous updates according to the threat group’s latest modus operandi. Also, threat actors may employ false flags by employing other groups of TTPs to disrupt the attribution method or hide their traces [21, 24].
@@
TABLE VIII: Attribution Methods and their Performance Comparison
@@
TABLE IX: Attribution Methods and their Performance Comparison
@@
    Close-world Solutions: APT groups evolve over time, and several threat groups emerge nowadays. It raises concerns regarding models trained on close-world solutions. Close-world solutions attribute threat groups to only what they have seen during training. The emergence of threat groups leads to misattribution when the given sample belongs to newly emerged groups whose sample has not been seen by the model during training. As a result, the trained model is attributed to a wrong threat group. Our review also finds that only [48] introduced the open-world solution, i.e., capability of their method for identifying unknown or newly emerged threat actors.
@@ VI Challenges
Attribution itself is a challenging and complex task and requires diverse consideration during the process. Additionally, APT attacks’ sophisticated, multifaceted, multi-staged, and dynamic nature makes it more difficult. The challenges faced by APT attributions are discussed in this section.
@@ VI-B False Flags
APT groups may deliberately leave false clues, known as false flags, to mislead investigators about their true identity [24, 10]. This deceptive tactic involves using tools, techniques, and methods typically associated with other groups or countries [21]. By adopting these recognizable signatures, they aim to create confusion and misdirection, making it appear that another entity is responsible for the attack. This strategy complicates the attribution process and increases the likelihood that investigators will draw incorrect conclusions about the origin and perpetrators of the cyber assault. False flags are the sophisticated method to obscure the true actors behind an attack, highlighting the complexity and challenges in accurate cyber threat attribution.
@@ VII-A How does one enhance the current automated APT attribution methodologies?
Cyber threat attribution significantly benefits from cross-disciplinary collaboration, combining insights from various fields such as cybersecurity, geopolitics, sociology, and intelligence analysis. This integrated approach ensures that technical data is enriched with a contextual understanding, providing a more comprehensive view of the threats. Researchers can explore automated frameworks and methodologies to foster such collaboration, creating systems where experts from different disciplines can contribute their unique perspectives. This interdisciplinary approach can lead to more robust and nuanced attribution conclusions, allowing for a deeper understanding of cyber attacks’ motives, methods, and implications. By leveraging the strengths of diverse fields, the attribution process can become more accurate and insightful, ultimately enhancing overall cybersecurity efforts.
@@ VII-F How can diverse data sources be effectively integrated and harmonized to improve the APT attribution?
Benchmark datasets and evaluation metrics are essential for advancing the field of cyber threat attribution. Creating standardized benchmark datasets enables researchers to assess and compare different attribution methods rigorously. These datasets provide a consistent basis for testing, ensuring that new techniques are evaluated against a common standard. Evaluation metrics, particularly those focused on precision, are crucial for measuring the accuracy and reliability of attribution methods. High precision ensures the methods correctly identify actual threat actors while minimizing false positives. Researchers can drive continuous improvement in attribution technologies by emphasizing precision in evaluation metrics, leading to more accurate and trustworthy results.
@@ VIII Conclusion
This survey provides an extensive review of the current landscape in cyber threat attribution and discusses the critical challenges and advancements that have shaped the field. The survey starts by exploring key artifacts that contribute to attribution and provides a taxonomy for scattered attribution artifacts. The taxonomy provides a standard way to understand and measure the possible attribution artifacts based on various qualitative measures. This survey also explores the available attribution datasets in the public domain and provides a classification of attribution datasets and their comparison. We further present the classification of available attribution methods based on the type of algorithm used and the dataset used for attribution. A comprehensive review of current literature on APT attribution is presented, along with critical comments and detailed discussions. Further, we analyze APT attribution’s current state and mention challenges toward advancing attribution methodologies. These comprehensive analysis of current literature led us to understand the future research directions in this domain and present the open research problems that researchers and practitioners can look into to advance the current state of APT attribution. By synthesizing these insights, this survey provides a comprehensive overview of the current state of APT attribution and offers valuable guidance for future research. We believe this work will inspire new advancements and practical solutions in threat attribution, ultimately fortifying our defenses against sophisticated cyber threats.
@@ Acknowledgments
[Artifact analysis reasoning] This section provides a logical analysis of the listed artifacts based on the selected measures.
TABLE X: Artifact analysis reasoning glimpse
@@
Indicator of Compromise (IoCs) Highly relevant as if IOCs like IP address, emails, cryptographic hashes are directly matched with any historical incident, then it may help in attribution confidently It can be easily manipulated, spoofed, or reused by different threat actors, reducing their integrity in providing reliable attribution evidence Often come from a mix of reliable open-source intelligence and private sector reports Available shortly after or even during an attack, allowing for quick detection and response Often easy to collect and analyze, as they are usually widely shared and well-documented
C&C Infrastructure Highly relevant as it often directly links to specific threat actors’ operations, providing a crucial link in identifying and attributing attacks. It can be easy for attackers to disguise or alter without compromising their operations Typically include in-depth network analysis and cybersecurity experts, making them highly credible and reliable for accurate attribution Can be identified and analyzed relatively quickly, but often requires ongoing monitoring and analysis to be effective in real-time Requires specific tools and expertise, and it can be challenging to detect active C&C servers
@@
Crypto Currency Transaction While they can be useful for tracking financial transactions, cryptocurrency transactions are often anonymized and lack direct relevance to specific threat actor attribution While the blockchain provides transparency, the anonymity and potential use of mixers reduce their reliability for attribution Often derived from blockchain analysis, which is transparent and traceable Often takes time due to the need to analyze blockchain data, and relevant information may only become clear after the transactions have been fully processed Tracking and analyzing cryptocurrency transactions can be complex due to the need for specialized tools and the anonymized nature of transactions
@@
Tactics, Techniques & Procedures (TTPs) Highly relevant as they represent the behavioral patterns of threat actors, which are often unique and difficult to fake, providing strong attribution evidence High integrity because they represent consistent behavioral patterns that are hard for attackers to alter without changing their operational methods Usually derived from detailed cybersecurity research and historical attack patterns, often published by reputable security organizations, ensuring high veracity Generally identified over time through pattern analysis and research, and often emerge after multiple incidents Accessible through threat intelligence reports and databases, but they require in-depth analysis and understanding to link them accurately to specific actors
@@
Malware & Code Analysis Analyzing malware and code can reveal unique signatures and methods that are often tied to specific threat groups, making this artifact highly relevant for attribution High integrity because unique code signatures and methods are difficult to fake or alter significantly without detection typically conducted by skilled cybersecurity researchers and forensic experts, ensuring that the sources are reliable and the findings are trustworthy Typically time-consuming, requiring in-depth investigation that can delay actionable insights Requires significant technical expertise and specialized tools to decompile and analyze code, making it less accessible to general analysts
@@
Toolchains The tools and software used by attackers can be indicative of specific groups, particularly when tied to known actor methodologies, making this highly relevant It can be changed based on attacker group’s preference and advancements in advanced tools Often linked to direct observations by security professionals, including forensic investigations, providing high veracity Requires detailed investigation, which can take time, but once identified, it can provide timely insights into ongoing campaigns Identifying and analyzing a threat actor’s toolchain requires detailed investigation and correlation of multiple artifacts, which can be resource-intensive
@@
Language & Writing Style Language and writing style can help identify specific individuals or groups, but they are less definitive on their own and may require corroboration with other artifacts While unique, it can be imitated or altered, requiring careful consideration to avoid misattribution Usually extracted from collected malware samples or received spam emails Typically takes time, especially if it requires gathering sufficient samples for a conclusive analysis, leading to low timeliness requires access to relevant samples and linguistic expertise, making it moderately accessible
@@
Public Claims and Hacker Forums While public claims and discussions on hacker forums can provide insights, they are not always reliable and may require validation, giving them medium relevance Low integrity due to the potential for false claims, misinformation, or deliberate deception by actors Often unverified and can be intentionally misleading, leading to a low level of veracity in these sources Can emerge quickly and sometimes delay relatively easy to access through open-source intelligence (OSINT), though its credibility varies
@@
Threat Intelligence Databases These databases compile various artifacts and intelligence but rely on the quality and context of the data, providing medium relevance in the attribution process It varies depending on the source and verification processes, making them moderately reliable for attribution Typically curated by established cybersecurity organizations, but sometimes may be updated by individual security researchers, which might raise concern regarding source integrity Updated regularly, but there can be a delay between the detection of new threats and the inclusion of information in the database These databases are designed to be accessible to analysts, providing structured information that is easy to query and analyze
Continue on the next page
TABLE XI: Artifact analysis reasoning glimpse
@@
Geopolitical It can offer valuable insights into potential motivations and likely actors, but it needs to be combined with other evidence for precise attribution Medium integrity, as it is interpretative and may be influenced by biases or incomplete information, but still provides valuable context Can be reliable, but they may also be influenced by political biases or incomplete information Requires a broad and often slow-gathering of information, making it less timely for immediate threat attribution and response Requires a good understanding of current geopolitical contexts and access to relevant intelligence reports
@@
Victimology Understanding the target’s profile can be highly relevant, as specific threat actors often target certain industries or regions, providing clues to their identity It relies on patterns that could potentially be aid if attackers intentionally target victims Often based on reported incidents, patterns and data collected from victims Can provide timely insights once sufficient data on targeted entities is gathered, but this often happens after multiple incidents have occurred Requires detailed analysis of attack patterns and affected entities, which can be resource-intensive but is generally accessible with proper data
@@
Post-Incident Communication It can offer insights into the actors involved, but it is less direct and may require interpretation, giving it medium relevance It can provide reliable insights, but may also be subject to manipulation or deception Usually comes from victim’s conversation with threat actor typically emerge with a delay This information can be accessible through monitoring communication channels, but it requires active intelligence gathering and analysis
@@
Time Zone Analysis While time zone analysis can suggest the likely origin of the threat actor, it is a circumstantial piece of evidence It is circumstantial and can be easily misrepresented by attackers using tools to obfuscate their actual location Often based on generally reliable metadata and logs can be conducted quickly using metadata from logs or other artifacts, providing timely insights into the possible origin of the attack Relatively straightforward and can often be conducted using metadata from logs and other time-stamped artifacts

### v3

@@ V-A Automated attribution framework
V-B APT attribution dataset and knowledge base classification
We curate the publicly available datasets and knowledge bases, which can be used to develop APT attribution methodology. In our exploration, we classify public data sources based on the type of data they consist. Therefore, we classify the data repositories into four types: Malware dataset, Threat report dataset, Attack pattern dataset, and Heterogeneous dataset. The distribution is shown in Fig. 6. We find a total of twelve public sources that facilitate data collection specific to APT attacks, which we discuss in detail below:
Figure 6: Dataset and knowledge base classification of APT attribution
@@ V-B2 Threat report dataset
V-B3 Attack pattern dataset
MITRE ATT&CK framework [118] is a knowledge base which lists a set of TTPs used by threat groups in past attributed attack campaigns. We can map the TTPs associated with a specific threat group through ATT&CK framework [80].
V-B4 Heterogeneous dataset
An Excel sheet focused on APT groups and their operations [103] is also a knowledge base and critical asset for cybersecurity analysts and researchers. It encompasses a wealth of data, including identifying APT groups, their geographical or organizational origins, details on TTPs employed, associated malware tools, descriptions of known campaigns, and IoCs. This organized compilation facilitates a comprehensive understanding of the threat landscape posed by these sophisticated actors, detailing their methodologies, and the specific vulnerabilities they exploit. By aggregating and categorizing this information, the document aids in identifying, analyzing, and mitigating threats from various APT groups, enhancing the overall cybersecurity posture against such entities.
MITRE’s knowledge base curates the extensive details related to activities performed by threat groups [106]. It endeavors to map APT groups to their TTPs [118], as well as the softwares [111] attackers utilize and the campaigns [115] they conduct, and various naming aliases which acknowledge that different organizations may track the same entities under different names. This approach aids in understanding the overlaps and distinctions among groups, enhancing analyst awareness and encouraging further investigation. Software in MITRE framework encompasses a variety of applications, ranging from custom malware to commonly used operating system utilities, categorized into malware and tools based on their intended use by adversaries. This categorization helps understand the specific techniques that different software instances can employ, mapped to the groups known to use them. Campaigns are defined as structured series of intrusion activities with specific objectives, targets, and duration. MITRE records these activities, providing unique identifiers for unnamed operations or using publicly reported names, which aims to attribute them to known groups and their used software where possible. MITRE facilitates a comprehensive view of threat actor’s operations over time and offers insights into the evolution of their tactics and methodologies.
Laurenza and Lazzeretti introduce “dAPTaset”  [26], a structured database for APT information. The database structure includes APTs, keywords, threat reports, samples, network information, CVEs (Common Vulnerabilities and Exposures), attack techniques, and more. The data collection process involves multiple sources such as MITRE ATT&CK [118], APT groups and operations spreadsheets [103], APTNotes [19], and MISP galaxy cluster [13], from which APT names, aliases, victims, operations, and external references are extracted. They parsed the public threat reports to gather IoCs, file hashes, IP addresses, network URLs, emails, CVEs, and keywords related to APTs. The dataset is further enriched by cross-referencing with cybersecurity websites like VirusTotal [16], MalwareBazaar [30], VirusShare [27], Malshare [34], and Malpedia [47] to validate and supplement information. This dataset uses $821$ reports of $88$ different threat groups. In addition, it also includes a total of $21,841$ binary hashes, $2,620$ network-related IoCs, $8,927$ unique files, $169$ CVEs, and $175$ different techniques related to APT activities. The dataset is publicly available [91] for the benefit of the community.
@@ V-B4 Heterogeneous dataset
V-C Classification of automated attribution methods
@@
V-C1 Malware-based methods
Kida et al. [29] present a novel way to leverage malware’s fuzzy hashes to attribute APT groups. They calculated five types of fuzzy hashes: TLSH, SSDEEP, SDHASH, IMPFUZZY, and LZJD for all malware samples and used computed hash as natural language input to the classifier. After hash calculation, the hashes are encoded using the NATO encoding method, and they use the padding method to keep the encoded length the same. Further, they use character level and n-gram vector embeddings to convert encoded hashes into feature vectors. The embedding vectors are forwarded to four different ML classifiers: RF, SVM, KNN, and NB. Their result demonstrates that RF with IMPFUZZY hash performs better than all implemented methods.
@@ V-C1 Malware-based methods
Rosenberg et al. [79] propose a method named DeepAPT to employ raw dynamic malware features and attribute APT groups linked with the given malware sample. They execute samples in a sandbox environment and collect behavior reports for each malware sample. DeepAPT considers words present in behavior reports as natural language features and evaluates the frequency of each unique word in all behavior reports. The authors selected the top $50,000$ frequent words as features and evaluated one-hot encoding to prepare the dataset. Further, they employ a 10-layer DNN to learn patterns from raw dynamic features for APT group classification.
@@
Haddadpajouh et al. [116] introduce a multi-view fuzzy consensus clustering model to attribute cyber threat payloads, specifically malware, to their respective actors. Their study follows a multi-view approach (multiple pieces of evidence) rather than single-view (single piece of evidence) analyses. The multi-view is generated using four categories of features: Header, System call, Byte code, and Opcode. The authors calculate multi-view for each feature category using binary vector (1 if present 0 if absent), count vector, frequency of occurrence, TF-IDF vector, and eigenvector of CFGs. Further, they utilize fuzzy pattern trees and multi-modal fuzzy classifiers for inference. Their research demonstrates that a multi-view approach is significantly better than single-view analyses for cyber threat attribution.
@@ V-C4 Heterogeneous data-based methods
Figure 7: Distribution of feature types (a), transformation methods (b), and algorithm types (c).
For malware-based methods, authors in literature extract the following types of features from malware samples: words from cuckoo sandbox reports, OpCode, ByteCode, API call, Header data, IR (Intermediate Representation) codes, strings, Fuzzy hashes (TLSH, SSDEEP, SDHASH, IMPFUZZY, and LZJD), CFG, and runtime modules. Threat report-based attribution methods extract words from threat intelligence reports as features. Attack pattern-based methods leverage TTPs as a feature for attribution. Meanwhile, heterogeneous data-based attribution methods leverage diverse sets of features like TTPs, software used, and heterogeneous attributed graph schemas. The distribution of feature types used in the literature is shown in Fig. 7(a). Out of all features mentioned in the literature TTPs and threat report texts has been used widely to present attribution method.
For feature transformation, the literature leverages the following different methods: One-Hot encoding, SMOBI word embedding, TF-IDF, Count, Statistical Characteristics, SIMVER word embedding, n-gram method, NATO encoding, Basic block characteristics, frequency analysis, Node embedding, BERT embedding, and Node2vec embedding. The distribution of type of transformation methods has been used in the literature is shown in Fig. 7(b). Out of all used transformation methods One-Hot encoding and TF-IDF has been used widely in the literature.
@@ V-D Discussion
(P1) As malware can be tweaked, distributed, or sold, i.e., given the rise of MaaS (Malware-as-a-Service) [33], it may lead to notable challenges in uncovering the actual origin of a cyber attack. Additionally, malwares’ behaviour-based threat attribution may provide an inconsistent view of threat actors when the same threat actor employs numerous malware in various campaigns. (P2) Malware-based threat attribution can be misleading in the case of collaborative and multi-layered attacks, as these attacks do not rely on a single instance of malware. (P3) For the threat-report-based attribution methods, authors employ threat report words as features for classification in the literature. Analysing the words of two threat reports shows the similarity between threat reports rather than the similarity between the threat actor’s behavior or modus operandi. Comparing words for document similarity has been studied widely in the natural language processing domain [28, 38]. (P4) Given P3, such methods need to reverse engineer and identify words contributing to the attribution or profiling of a specific threat actor. (P5) Attack pattern-based methods leverage TTPs in the literature. TTP of a threat group can evolve over time [52], and TTP-based threat attribution can miss-attribute in such cases. Therefore, TTP-based attribution methods demand continuous updates according to the threat group’s latest modus operandi. Also, threat actors may employ false flags by employing other groups of TTPs to disrupt the attribution method or hide their traces [21, 24].
@@
TABLE VIII: Comparison of attribution methods and their performance
@@
TABLE IX: Comparison of attribution methods and their performance
@@
    Lacks open-world solutions: APT groups evolve over time, and several threat groups emerge nowadays. It raises concerns regarding models trained on close-world solutions. Close-world solutions attribute threat groups to only what they have seen during training. The emergence of threat groups leads to misattribution when the given sample belongs to newly emerged groups whose sample has not been seen by the model during training. As a result, the trained model is attributed to a wrong threat group. Our review also finds that only [48] introduced the open-world solution, i.e., capability of their method for identifying unknown or newly emerged threat actors.
@@ VI Challenges
Attribution itself is a challenging and complex task and requires diverse consideration during the process. Additionally, APT attacks’ sophisticated, multifaceted, multi-staged, and dynamic nature makes it more difficult. The challenges observed by APT attribution methods are discussed in this section.
@@ VI-B False flags
APT groups may deliberately leave false clues, known as false flags, to mislead investigators about their true identity [24, 10, 4]. This deceptive tactic involves using tools, techniques, and methods typically associated with other groups or countries [21, 104]. By adopting these recognizable signatures, they aim to create confusion and misdirection, making it appear that another entity is responsible for the attack. This strategy complicates the attribution process and increases the likelihood that investigators will draw incorrect conclusions about the origin and perpetrators of the cyber assault. False flags are the sophisticated method to obscure the true actors behind an attack, highlighting the complexity and challenges in accurate cyber threat attribution.
@@ VII-A How does one enhance the current automated APT attribution methodologies?
Cyber threat attribution significantly benefits from cross-disciplinary collaboration, combining insights from various fields such as cybersecurity, geopolitics, and open-source and signal intelligence analysis. This integrated approach ensures that technical data is enriched with a contextual understanding, providing a more comprehensive view of the threats. Researchers can explore automated frameworks and methodologies to foster such collaboration, creating systems where experts from different disciplines can contribute their unique perspectives. This interdisciplinary approach can lead to more robust and nuanced attribution conclusions, allowing for a deeper understanding of cyber attacks’ motives, methods, and implications. By leveraging the strengths of diverse fields, the attribution process can become more accurate and insightful, ultimately enhancing overall cybersecurity efforts.
@@ VII-F How can diverse data sources be effectively integrated and harmonized to improve the APT attribution?
Benchmark datasets and evaluation metrics are essential for advancing the field of cyber threat attribution. Creating standardized benchmark datasets enable researchers to assess and compare different attribution methods rigorously. These datasets provide a consistent basis for testing, ensuring that new techniques are evaluated against a common standard. Evaluation metrics, particularly those focused on precision, are crucial for measuring the accuracy and reliability of attribution methods. High precision ensures the methods correctly identify actual threat actors while minimizing false positives. Researchers can drive continuous improvement in attribution technologies by emphasizing precision in evaluation metrics, leading to more accurate and trustworthy results.
@@ VIII Conclusion
This survey provides an extensive review of the current landscape in cyber threat attribution and discusses the critical challenges and advancements that have shaped the field. The survey starts by exploring key artifacts that contribute to attribution and provides a taxonomy for scattered attribution artifacts. The taxonomy provides a standard way to understand and measure the possible attribution artifacts based on various qualitative measures. This survey also explores the available attribution datasets in the public domain and provides a classification of attribution datasets and their comparison. We further present the classification of available attribution methods based on the type of algorithm used and the dataset used for attribution. A comprehensive review of current literature on automated APT attribution is presented, along with critical comments and detailed discussions. Further, we analyze APT attribution’s current state and mention challenges toward advancing attribution methodologies. These comprehensive analysis of current literature led us to understand the future research directions in this domain and present the open research problems that researchers and practitioners can look into to advance the current state of APT attribution. By synthesizing these insights, this survey provides a comprehensive overview of the current state of APT attribution and offers valuable guidance for future research. We believe this work will inspire new advancements and practical solutions in threat attribution, ultimately fortifying our defenses against sophisticated cyber threats.
@@ Acknowledgments
[Artifact analysis reasoning] This section presents a rationale for analyzing the listed artifacts, based on the chosen evaluation measure.
TABLE X: Artifact analysis rationale glimpse
@@
Indicator of Compromise (IoCs) IoCs like IP address, emails can be directly matched with any historical incident Can be easily manipulated, spoofed, or reused by different threat actors Often come from various reliable open-source intelligence and private sector reports Available shortly after or even during an attack, allowing for quick detection and response Often easy to collect and analyze, as they are usually widely shared and well-documented
C&C Infrastructure Evidence of using same infrastructure such as C&C server can provide crucial link Easy for attackers to disguise or alter without compromising their operations Analyzed from network traffic analysis and collected from reliable open-source intelligence Can be identified and analyzed relatively quickly, but often requires ongoing monitoring and analysis Requires specific tools and third-party database, and it can be challenging to analyze active C&C servers
@@
Cryptocurrency Transaction These transactions are often anonymized and lack direct relevance to specific threat actor The anonymity and potential use of mixers reduce reliability for attribution Often derived from blockchain analysis, which is transparent and traceable Often takes time due to the need to analyze blockchain data, and relevant information may only become clear after the transactions have been fully processed Tracking and analyzing these transactions can be complex due to the need for specialized tools and the anonymized nature of transactions
@@
Tactics, Techniques & Procedures (TTPs) Represent the behavioral patterns of threat actors, which are often unique and difficult to fake According to Pyramid of Pain, it is hard for attackers to alter without changing their operational methods Derived from detailed cybersecurity research and historical attack patterns, often published by reputable security firms Generally identified through pattern analysis and research, and often emerge after analyzing multiple incidents Accessible through threat intelligence reports and databases, but require in-depth analysis and understanding to link specific threat actor
@@
Malware & Code Analysis Reveal unique signatures and methods that are often tied to specific threat groups Unique code signatures and methods are difficult to fake or alter significantly without detection Conducted by skilled cybersecurity researchers and forensic experts, ensuring that the findings are trustworthy Can be time-consuming, requiring in-depth investigation and analysis to uncover full spectrum of maliciousness Requires significant technical expertise and specialized tools to decompile and analyze code
@@
Toolchains Can be indicative of specific groups, particularly when tied to known actor methodologies Can be changed based on attacker group’s preference and advancements in advanced tools Often linked to direct observations by security professionals, including forensic investigations, providing high veracity Requires detailed investigation, which can take time, but once identified, it can provide timely insights into ongoing campaigns Identifying and analyzing a threat actor’s toolchain requires detailed investigation and correlation of multiple artifacts
@@
Language & Writing Style Less definitive on their own and may require corroboration with other artifacts While unique, it can be imitated or altered, requiring careful consideration Can be easily extracted from collected malware samples or received spam emails Typically takes time, especially if it requires gathering sufficient samples for a conclusive analysis Requires access to relevant samples and linguistic expertise
@@
Public Claims and Hacker Forums Not always relevant and reliable, may require validation Potential for false claims, misinformation, or deliberate deception by actors Often unverified and can be intentionally misleading Can emerge quickly and sometimes delay depends on threat actor modus operandi Relatively easy to access through OSINT, though its credibility varies
@@
Threat Intelligence Databases Compile various artifacts and intelligence but rely on the quality and context of the data Varies depending on the source and verification processes, making them moderately reliable Curated by established cybersecurity firms, but can be updated by non-credible individuals Updated regularly, but there can be a delay between the detection of new threats and the information inclusion in the database Designed to be easily accessible to analysts, providing structured information that is easy to query and analyze
@@
Geopolitical Can offer valuable insights into potential motivations and likely actors, but needs to be combined with other relevant evidences Interpretative but may be influenced by biases or incomplete information, still provides valuable context in some cases Can be reliable, but may also be influenced by political biases or incomplete information Requires a broad and often slow-gathering of information Requires a good understanding of current geopolitical contexts and access to relevant intelligence reports
@@
Victimology Highly relevant, as specific threat actors often target certain industries or regions, providing clues to their identity Relies on patterns that could potentially be aid if attackers intentionally target victims Often based on reported incidents, patterns and data collected from victims Can provide timely insights once sufficient data on targeted entities is gathered Requires detailed analysis of attack patterns and affected entities, which can be resource-intensive
@@
Post-Incident Communication May offer insights into the actors involved, but it is less direct and may require interpretation May provide reliable insights, but may also be subject to manipulation or deception Usually comes from victim’s conversation with threat actor Typically emerge with a delay in engaging threat actor in communication Can be accessible through monitoring communication channels, but requires active intelligence gathering
@@
Time Zone Analysis Can suggest the likely origin of the threat actor, it is a circumstantial piece of evidence Circumstantial and can be easily misrepresented by attackers using tools to obfuscate their actual location Often based on generally reliable metadata and logs Conducted quickly by leveraging metadata from logs or other artifacts, providing timely insights into potential origins Relatively straightforward and can often be accessible from metadata of logs and other time-stamped artifacts